# HEIST: A Graph Foundation Model for Spatial Transcriptomics and Proteomics Data

### v2

@@ Appendix B Data Preprocessing and Graph creation
Data Preprocessing. We begin by importing raw spatial transcriptomics data, followed by standard preprocessing steps, including outlier removal, gene expression normalization, and gene filtering to retain highly variable genes. To mitigate technical noise and dropout effects common in single-cell measurements, we apply MAGIC (Dijk et al., 2017) to denoise gene expression data before all downstream computations.
Cell-Cell Graph Construction. To capture spatial relationships between cells, we construct a cell-cell graph based on physical proximity. Spatial locations are used to compute Voronoi polygons, which define local neighborhoods. Edges in the graph are assigned based on adjacency in the Voronoi diagram, ensuring that the graph structure accurately reflects tissue architecture and local microenvironments.
Cell-Type-Specific Gene Co-Expression Networks Construction. For each cell type, we build a gene co-expression network that captures functional dependencies between genes. If cell type annotations are provided, we use them directly; otherwise, we infer cell types via Leiden clustering on the denoised gene expression data. After cell types are assigned, we compute pairwise mutual information (MI) between genes within each cell type using the denoised expression values. Gene pairs with MI greater than a threshold $\tau$ are connected by an edge in the gene co-expression network, resulting in a sparse, cell-type-specific co-expression graph. This process captures gene-gene co-expression patterns and co-expression dependencies unique to each cellular context.
@@ Appendix C Implementation details
In this section, we explain the implementation details including the positional encoding initialization, graph transformer architecture, and hyperparameters used in pre-training phase.
@@ C.1 Positional encodings
HEIST incorporates positional encodings (PE) at both the cell and gene levels to inject spatial and co-expression structure into the learned representations. As opposed to traditional graph PEs such has Laplacian PE (Dwivedi & Bresson, 2020), random walk PE (Dwivedi & Bresson, 2020), node-centrality based PE (Ying et al., 2021), we make use of sinusoidal PEs. This design choice is motivated by the fact that sinusoidal PE yields expressive representations and are computationally in-expensive as opposed to traditional graph PEs, leading to efficient and expressive representations (Wen et al., 2023b). Below, we describe how these encodings are constructed and why they are suitable for spatial transcriptomics.
Each cell is associated with a spatial coordinate $(x,y)\in\mathbb{R}^{2}$, corresponding to its location in the tissue. To encode spatial information, we apply a two-dimensional extension of the sinusoidal positional encoding introduced in transformers (Vaswani et al., 2017):
@@
$\text{PE}_{c}=\texttt{2DSinusoidal}({\mathbf{P}}),\text{PE}^{k}_{g}=\texttt{RankSinusoidal}({\mathbf{X}}_{k});\quad{\mathbf{H}}_{c}^{(0)}=\text{PE}_{c}+{\mathbf{P}},{\mathbf{H}}_{g}^{k^{(0)}}=\text{PE}^{k}_{g}+{\mathbf{X}}_{k}$
where ${\mathbf{H}}_{c}^{(0)}$ and ${\mathbf{H}}_{g}^{k^{(0)}}$ are the input cell and gene embeddings for cell $k$, respectively. Together, the PE helps HEIST incorporate spatial geometry and transcriptional ordering in a biologically meaningful and computation-efficient way.
@@ C.2 Intra-level message passing
We incorporate biological inductive biases by applying global (all-to-all) attention on the cell graph to capture long-range spatial interactions and local attention that depends on the graph structure of gene graphs to model fine-grained co-expression patterns (Papalexi & Satija, 2018). Capturing global dependencies is crucial for detecting large-scale tissue structures, such as immune infiltration zones and tumor-stroma boundaries (Zhu et al., 2024), while modeling local co-expression networks enables identification of tightly coordinated gene modules within specific niches (Halpern et al., 2017). HEIST performs intra-level message passing independently at the cell and gene levels to capture both local and global dependencies within each modality.
@@
where ${\mathbf{H}}_{c}^{(l)}$ denotes the cell embeddings at layer $l$, and ${\mathbf{A}}_{c}$ is the adjacency matrix of the spatial cell graph. The transformer layer enables long-range interactions across spatial regions, while the GIN layer aggregates local neighborhood information. This hybrid approach captures both global spatial context (e.g., tissue-level structure) and localized microenvironments (e.g., niches or boundaries).
@@ C.3 Hyperparameters
Activation function GeLu (Hendrycks & Gimpel, 2016)
Optimizer AdamW (Loshchilov & Hutter, 2017)
@@ D.1 Pretraining dataset
In contrast to previous models that incorporate lower-resolution data from Visium arrays (Cui et al., 2024), HEIST focuses exclusively on single-cell resolution datasets. This design choice reflects the fact that Visium data lacks true single-cell resolution, instead capturing transcriptomic signals aggregated over spatial spots that often contain heterogeneous mixtures of multiple cell types. Such coarse-resolution data can introduce confounding signals and limit the model’s ability to accurately learn cell-level spatial dependencies and gene co-expression. By focusing on single-cell spatial transcriptomics, HEIST is better positioned to capture fine-grained spatial organization, cell-cell interactions, and context-dependent gene co-expreesion critical for modeling tissue microenvironments and cellular heterogeneity.
@@ D.2 Downstream Tasks
Tissue-level classification. In this task, the goal is to predict tissue-level outcomes, including response to immunotherapy, treatment outcomes, remission status, and placenta condition. Accurate classification requires capturing both local cellular microenvironments and broader tissue organization patterns. We first evaluate HEIST on datasets from Charville (Colon), UPMC (Neck), and DFCI (Neck), collected using CO-Detection by Indexing (CODEX) technology. We also predict immunotherapy response on a skin cancer dataset (Ptacek et al., 2021) collected through multiplex ion beam imaging (MIBI). Finally, we classify placenta condition into normal placenta, placenta accreta spectrum (PAS), and placental insufficiency using data collected with Xenium. Spatial modeling is crucial for placenta classification, as conditions like PAS involve disrupted tissue architecture and abnormal cell invasion, which gene expression alone cannot fully capture. For this task, we extract cell representations from a frozen HEIST encoder and train an MLP head. We report AUC-ROC as the evaluation metric.
Gene imputation. In spatial transcriptomics, gene expression profiles are often incomplete or noisy due to technical limitations and measurement dropouts. Gene imputation aims to predict missing gene expression values, recovering biologically plausible expression patterns. HEIST performs this task by leveraging the learned gene embeddings, which capture both co-expression relationships from gene co-expression networks and spatial dependencies from their parent cells. Through cross-level message passing, the model can incorporate spatial cues to refine gene predictions, enabling more accurate reconstruction of missing data. We select the genes to be masked using stratified sampling according to gene sparsity (Avşar & Pir, 2023) and mask these genes during the fine-tuning phase. This task can be performed in a zero-shot fashion or by fine-tuning the HEIST decoder on the downstream task. We report Pearson correlation as the evaluation metric for this task.
@@ D.3 Downstream datasets
We have curated different datasets for each downstream prediction task. To show our models generalize well we specifically used 4 different technologies across the different tasks. The dataset description can be found on Table A2
@@
Appendix E Extra results
@@
Figure A2: Comparisons between embeddings focusing on L4-IT neurons. The figure shows the cell type embeddings, followed by spectral clustering and a plot of all clusters that contain the specific cell type. A. scGPT-spatial results. B. CellPLM results C. HEIST results (same as Figure 3) D. Ground Truth cell types colored in space.
@@
Runtime comparison. Table A4 compares the runtime efficiency of HEIST with existing foundation models for generating embeddings from a tissue sample with 19,826 cells. HEIST achieves substantial speed improvements, being 8× faster than scGPT and 48× faster than scFoundation, while providing both cell and gene embeddings. Although CellPLM is faster, it does not compute gene representations, making it less suitable for tasks requiring joint cell and gene analysis.
Table A4: Runtime (in seconds) comparison to extract embeddings from a tissue with 19826 cells. * CellPLM does not compute gene representations.
Model Runtime
scFoundation 290.48
scGPT 49.37
HEIST 6.69
CellPLM* 1.97

### v3

@@ Appendix B Data Preprocessing and Graph creation
Data Preprocessing. We begin by importing raw spatial transcriptomics data, followed by standard preprocessing steps, including outlier removal, gene expression normalization, and gene filtering to retain highly variable genes. We denote the raw matrix by $X_{\mathrm{raw}}\in\mathbb{R}^{N\times G}$ and the normalized matrix by $X$, with $X_{ij}=\mathrm{Norm}(X_{\mathrm{raw},ij}).$ To mitigate technical noise and dropout effects common in single-cell measurements, we apply MAGIC (dijk2017magic) to denoise gene expression data before all downstream computations. Formally,
$\tilde{X}=\mathrm{MAGIC}(X).$
Cell-Cell Graph Construction. To capture spatial relationships between cells, we construct a cell-cell graph based on physical proximity. Each cell has spatial coordinates $p_{i}=(x_{i},y_{i})\in\mathbb{R}^{2}$. Spatial locations are used to compute Voronoi polygons, which define local neighborhoods. Edges in the graph are assigned based on adjacency in the Voronoi diagram, ensuring that the graph structure accurately reflects tissue architecture and local microenvironments. We write this as
$(i,j)\in\mathcal{E}_{c}\;\Longleftrightarrow\;\mathrm{Vor}(p_{i})\ \text{and}\ \mathrm{Vor}(p_{j})\ \text{share a boundary}.$
[…]
After cell types are assigned, we compute pairwise mutual information (MI) between genes within each cell type using the denoised expression values. Gene pairs with MI greater than a threshold $\tau$ are connected by an edge in the gene co-expression network, resulting in a sparse, cell-type-specific co-expression graph. This process captures gene-gene co-expression patterns and co-expression dependencies unique to each cellular context. For a cell type $t_{k}$ with cell index set $\mathcal{I}_{k}=\{i:t_{i}=t_{k}\}$,
$\mathrm{MI}_{ab}^{(k)}=I\left(\tilde{X}^{(k)}_{\cdot,a};\tilde{X}^{(k)}_{\cdot,b}\right),\qquad(a,b)\in\mathcal{E}_{g}^{(k)}\Longleftrightarrow\mathrm{MI}_{ab}^{(k)}\geq\tau_{k}.$
Choice of $\tau$. The threshold $\tau$ used for constructing the gene co-expression networks is data-driven and cell-type-specific. Specifically,
$\tau_{k}=\mu_{k}+\sigma_{k},\qquad\mu_{k}=\mathbb{E}_{a\neq b}[\mathrm{MI}_{ab}^{(k)}],\qquad\sigma_{k}=\mathrm{Std}\big(\{\mathrm{MI}_{ab}^{(k)}\}_{a\neq b}\big).$
This adaptive choice yields graphs with  20–30% edge density, which empirically provides sufficient structure to capture co-expression dynamics without over-saturating the graph. In earlier development, we experimented with a constant global $\tau$, but this produced artifacts: some cell types became overly dense while others became nearly edge-free. This mismatch harms hierarchical message passing and leads to degraded downstream performance. We also noticed similar trends in graph based baselines such as STAGATE, GraphST, and Novae, where we had to tune the thresholds manually to control the connectivity. We have highlighted these details in Appendix Section B.
Batching the tissue. Each tissue is divided into spatially coherent blocks that contain only a subset of nearby cells. Let $B$ denote the block size and $\{\mathcal{B}_{1},\dots,\mathcal{B}_{M}\}$ be the resulting partition with $|\mathcal{B}_{m}|\leq B$. In practice, we choose the largest block size that fits within the memory budget of a single GPU; for our hardware (40GB L40), this is 256 cells per block. Larger blocks allow HEIST to capture mid-range spatial interactions more effectively, while still avoiding the quadratic cost of full-tissue attention. We expect diminishing returns beyond a certain block size: once the local microenvironment and its spatial context are fully captured, adding more distant cells contributes little additional signal.
@@
Final Output. The final preprocessing pipeline produces:
  • A spatial $\mathcal{G}_{c}(\mathcal{C},\mathcal{E}_{c},P)$ cell-cell graph encoding local tissue structure.
  • Cell-type-specific gene co-expression networks $\{\mathcal{G}_{g}^{t_{k}}(\mathcal{V},\mathcal{E}_{t_{k}},X_{k})\}_{k=0}^{|\mathcal{C}|}$ capturing co-expression function within each cell type.
@@ Appendix C Implementation details
In this section, we explain the implementation details including batching, the positional encoding initialization, graph transformer architecture, and hyperparameters used in pre-training phase.
@@ C.1 Positional encodings
HEIST incorporates positional encodings (PE) at both the cell and gene levels to inject spatial and co-expression structure into the learned representations. As opposed to traditional graph PEs such has Laplacian PE (dwivedi2020generalization), random walk PE (dwivedi2020generalization), node-centrality based PE (ying2021transformers), we make use of sinusoidal PEs. This design choice is motivated by the fact that sinusoidal PE yields expressive representations and are computationally in-expensive as opposed to traditional graph PEs, leading to efficient and expressive representations (wen2023single). Below, we describe how these encodings are constructed and why they are suitable for spatial transcriptomics.
Each cell is associated with a spatial coordinate $(x,y)\in\mathbb{R}^{2}$, corresponding to its location in the tissue. To encode spatial information, we apply a two-dimensional extension of the sinusoidal positional encoding introduced in transformers (vaswani2017attention):
@@
$\text{PE}_{c}=\texttt{2DSinusoidal}({\mathbf{P}}),\text{PE}^{k}_{g}=\texttt{RankSinusoidal}({\mathbf{X}}_{k});\quad{\mathbf{H}}_{c}^{(0)}=\text{PE}_{c}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbf{P}}_{c}^{\rm repeat}},{\mathbf{H}}_{g}^{k^{(0)}}=\text{PE}^{k}_{g}+{\mathbf{X}}_{k}$
[…]
@@ C.2 Intra-level message passing
t Refer to caption]
Figure A2: HEISTLayer Block
We incorporate biological inductive biases by applying global (all-to-all) attention on the cell graph to capture long-range spatial interactions and local attention that depends on the graph structure of gene graphs to model fine-grained co-expression patterns (papalexi2018single). Capturing global dependencies is crucial for detecting large-scale tissue structures, such as immune infiltration zones and tumor-stroma boundaries (zhu2024mapping), while modeling local co-expression networks enables identification of tightly coordinated gene modules within specific niches (halpern2017single). HEIST performs intra-level message passing independently at the cell and gene levels to capture both local and global dependencies within each modality.
@@
where ${\mathbf{H}}_{c}^{(l)}$ denotes the cell embeddings at layer $l$, and ${\mathbf{A}}_{c}$ is the adjacency matrix of the spatial cell graph. The transformer layer enables long-range interactions across spatial regions, while the GIN layer aggregates local neighborhood information. This hybrid approach captures both global spatial context (e.g., tissue-level structure) and localized microenvironments (e.g., niches or boundaries). Although the attention module is fully dense within each block, the computational cost is significantly reduced because the effective sequence length becomes the block size rather than the total number of cells in the tissue. A naive full-tissue attention operation would scale as $O(c^{2})$, which is expensive for large tissues. After spatial batching, the complexity becomes $\mathcal{O}(b^{2})$, where $b$ is the block size. In practice, $b$ is much smaller than the total number of cells in a tissue, so this reduction makes attention computation tractable while still capturing local and mid-range spatial interactions.
@@ C.3 Hyperparameters
Activation function GeLu (gelu)
Optimizer AdamW (loshchilov2017decoupled)
@@ D.1 Pretraining dataset
In contrast to previous models that incorporate lower-resolution data from Visium arrays (cui2024scgpt), HEIST focuses exclusively on single-cell resolution datasets. This design choice reflects the fact that Visium data lacks true single-cell resolution, instead capturing transcriptomic signals aggregated over spatial spots that often contain heterogeneous mixtures of multiple cell types. Such coarse-resolution data can introduce confounding signals and limit the model’s ability to accurately learn cell-level spatial dependencies and gene co-expression. By focusing on single-cell spatial transcriptomics, HEIST is better positioned to capture fine-grained spatial organization, cell-cell interactions, and context-dependent gene co-expreesion critical for modeling tissue microenvironments and cellular heterogeneity.
@@ D.2 Downstream Tasks
Tissue-level classification. In this task, the goal is to predict tissue-level outcomes, including response to immunotherapy, treatment outcomes, remission status, and placenta condition. Accurate classification requires capturing both local cellular microenvironments and broader tissue organization patterns. We first evaluate HEIST on datasets from Charville (Colon), UPMC (Neck), and DFCI (Neck), collected using CO-Detection by Indexing (CODEX) technology. We also predict immunotherapy response on a skin cancer dataset (melanoma-data) collected through multiplex ion beam imaging (MIBI). Finally, we classify placenta condition into normal placenta, placenta accreta spectrum (PAS), and placental insufficiency using data collected with Xenium. Spatial modeling is crucial for placenta classification, as conditions like PAS involve disrupted tissue architecture and abnormal cell invasion, which gene expression alone cannot fully capture. For this task, we extract cell representations from a frozen HEIST encoder, we aggregate the cell embeddings by taking a mean, and train an MLP head. We report AUC-ROC as the evaluation metric.
Gene imputation. In spatial transcriptomics, gene expression profiles are often incomplete or noisy due to technical limitations and measurement dropouts. Gene imputation aims to predict missing gene expression values, recovering biologically plausible expression patterns. HEIST performs this task by leveraging the learned gene embeddings, which capture both co-expression relationships from gene co-expression networks and spatial dependencies from their parent cells. Through cross-level message passing, the model can incorporate spatial cues to refine gene predictions, enabling more accurate reconstruction of missing data. We select the genes to be masked using stratified sampling according to gene sparsity (avcsar2023comparative) and mask these genes during the fine-tuning phase. For the skin dataset, which contains 30 proteomic markers, we masked 4 genes using stratified sampling based on gene sparsity following Avsar et al., corresponding to a  14% masking fraction. For the placenta dataset, we masked 16 out of 200 genes ( 8% masking). This task can be performed in a zero-shot fashion or by fine-tuning the HEIST decoder on the downstream task. We report Pearson correlation as the evaluation metric for this task.
@@ D.3 Downstream datasets
We have curated different datasets for each downstream prediction task. To show our models generalize well we specifically used 4 different technologies across the different tasks. The dataset description can be found on Table A2.
@@
Appendix E Additional results
@@
Figure A3: Comparisons between embeddings focusing on L4-IT neurons. The figure shows the cell type embeddings, followed by spectral clustering and a plot of all clusters that contain the specific cell type. A. scGPT-spatial results. B. CellPLM results C. STAGATE results D. GraphST results E. HEIST results (same as Figure 3) F. Ground Truth cell types colored in space.
@@
Runtime comparison. Table A4 compares the runtime efficiency of HEIST with existing foundation models for generating embeddings from a tissue sample with 19,826 cells. HEIST achieves substantial speed improvements, being 8× faster than scGPT and 48× faster than scFoundation, while providing both cell and gene embeddings. We also present show the coputational complexity analysis of existing models. We note that HEIST is linear in gene edges and gene count, one HEISTLayer is faster than one scGPT-spatial and scFoundation layer, however it is slower than CellPLM. Although CellPLM is faster, it does not compute gene representations, making it less suitable for tasks requiring joint cell and gene analysis.
Table A4: Runtime comparison for embedding extraction from a tissue with 19826 cells, and computational complexity of each model. *CellPLM does not compute gene representations.
Model Runtime (s) Complexity Interpretation
scFoundation 290.48 $\mathcal{O}(C^{2}G_{\text{nz}}^{2}d+CTd)$ Quadratic in cell count and nonzero genes. Decoder is linear due to Performer attention.
scGPT 49.37 $\mathcal{O}(T^{2}d+Td^{2})$ Quadratic in cell times number of gene tokens per spot
HEIST 6.69 $\mathcal{O}\big(d(C^{2}+E_{c}+CE_{g}+CT)\big)$ Quadratic in cells (global spatial transformer), linear in gene edges and gene count.
CellPLM* 1.97 $\mathcal{O}(C^{2}d)$ Quadratic in cells, linear in hidden size. No dependence on number of genes after aggregation.
@@
$\tau$ choice ablation. In Table A6, we compare the results of tissue classification between adaptive and constant threshold to create gene co-expression graphs. As we can see, the adaptive $\tau$ results consistently outperforms the constant version, capturing fine-grained co-expression structure.
Table A6: Ablation comparing constant and adaptive choice of $\tau$.
Dataset Charville Charville UPMC UPMC DFCI
$\tau$ Task Outcome Recurrence Outcome Recurrence Outcome
Constant 0.68 0.70 0.828 0.75 0.916
Adaptive 0.86 0.88 0.835 0.92 0.937
Importance of contrastive terms. To understand whether the three contrastive components interact positively, we carried out an ablation experiment in which we removed each term in turn: cell to cell, gene to gene, and cell to gene. As shown in Table A7, removing any single term consistently reduces performance across all datasets. Removing the cell to cell objective reduces Charville outcome prediction from 0.861 to 0.380. Removing the gene to gene objective produces the largest reduction on SEA cell classification, which drops from 0.995 to 0.610. These results show that each contrastive component captures complementary structural signals that include local cell similarity, gene program structure, and cross scale coupling between cells and genes. Using all three terms together provides the strongest supervision signal and does not create negative interactions. We also tested with a joint contrastive loss function that directly models the c–c, g–g, and c–g correlations together rather than simply adding the three contrastive losses. As we can see, the performance with this joint contrastive loss decreased slightly. The primary reason is that the joint contrastive objective significantly increased memory consumption, causing repeated OOM errors. To stabilize training, we were forced to reduce the batch size. Smaller batches provide far less microenvironmental diversity within each step, which is essential for learning strong contrastive signals. This reduced context ultimately impaired the effectiveness of the joint contrastive loss and led to the observed decrease in performance.
Table A7: Ablation of HEIST’s contrastive loss components across three tasks. Removing any of the three contrastive terms degrades performance across modalities, confirming that the full joint loss is necessary.
Loss Charville-Outcome Skin-Imputation SEA-Cell classification
Full contrastive $0.861\pm 0.086$ $0.807\pm 0.020$ $0.995\pm 0.015$
Joint contrastive 0.808 $\pm$ 0.040 0.775 $\pm$ 0.017 0.994 $\pm$ 0.004
No $c\leftrightarrow c$ $0.380\pm 0.119$ $0.570\pm 0.060$ $0.993\pm 0.006$
No $g\leftrightarrow g$ $0.575\pm 0.049$ $0.477\pm 0.033$ $0.610\pm 0.097$
No $g\leftrightarrow c$ $0.559\pm 0.065$ $0.693\pm 0.057$ $0.928\pm 0.002$
Attention interpretation. We analyzed the cell to cell attention scores learned by HEIST and CellPLM and visualized their top attention edges. We aggregated the attention across layers and heads to get a single attention maps. We only compared with CellPLM since that is the only other foundation model that calculates cell-cell attention.
As shown in Figure A4, the edges with the highest attention values in HEIST form clear, spatially coherent microenvironments. We observe groups of cells of different types connected in structured patterns, often resembling local niches such. These visual patterns indicate that HEIST’s attention mechanism focuses on meaningful neighborhoods rather than isolated pairs. In contrast, the top-attention edges from CellPLM are diffuse and scattered across the tissue. The highlighted edges do not show consistent spatial structure or cell-type organization. This suggests that CellPLM attention is much noisier and does not preferentially emphasize biologically relevant areas.
To quantify whether these visual differences reflect underlying biological communication, we compared all edges to curated ligand-receptor (LR) interactions from CellPhoneDB. Across the entire dataset, approximately 16 percent of edges correspond to known LR pairs. Randomly sampling 1000 edges also gives about 16 percent, which serves as the expected background rate. As shown in Table A8
HEIST shows strong enrichment for LR edges:
  • The top 100 high-attention edges contain 31 percent LR interactions.
  • The top 1000 edges contain 27.9 percent LR interactions.
  • This is better than chance which yields about 16 percent expected LR pairs.
CellPLM shows the opposite trend. Only 11 percent of its top 100 edges and 8.8 percent of its top 1000 edges correspond to LR interactions. This is clearly below the background rate, which indicates that CellPLM attention values do not align with known biological communication.
Figure A4: Visualizing edges with highest cell-cell attentions. We can clearly see that HEIST is able to form niches, while CellPLM high attention is scattered around the tissue.
Table A8: LR-rate measured using cell-cell attention.
Model LR pairs in top 100 LR pairs in top 1000
Background (all edges) 0.165 0.165
Random 0.164 0.165
CellPLM 0.110 0.088
HEIST 0.310 0.279
Ligand-receptor pair prediction. We train a linear probe to predict ligand–receptor (LR) interactions for all edges in the tissue graphs. For each model, we extract the cell embeddings, concatenate the embeddings of the two cells involved in a pair, and pass this concatenated representation through a simple single layer perceptron that predicts whether the pair corresponds to an LR interaction. As shown in the Table A9, HEIST achieves the highest AUC-ROC and outperforms all baselines by a clear margin. In particular, HEIST reaches an AUC-ROC of 0.995 $\pm$ 0.002. This result combined with the attention based results indicates that its embeddings capture the spatial signatures of LR communication more effectively than scFoundation, CellPLM, and scGPT-spatial.
Table A9: Predicting LR pairs
Model AUC-ROC
scFoundation $0.963\pm 0.004$
CellPLM $0.930\pm 0.006$
scGPT-spatial $0.984\pm 0.003$
HEIST $\mathbf{0.995\pm 0.002}$
Recovering unique niches based on tissue conditions. We compare the attention-derived niches across Braak stages, a clear progression emerges in Figure A5. In Braak 0, the high-scoring edges form small, isolated neuronal clusters that are relatively sparse and spatially compact. By Braak IV and V, these clusters become more numerous and begin to involve a broader set of cell types, indicating early signs of local microenvironmental restructuring. In Braak VI, the niches become much denser and more spatially extended. This gradual transition from isolated neuronal interactions to larger mixed-cell hubs suggests that HEIST attention captures the increasing microenvironmental complexity and tissue reorganization that accompany the progression of pathology across Braak stages.
Figure A5: HEIST attention reveals unique microenvironments depending on Braak stages
PE initialization. Our choice to incorporate spatial coordinates and gene-expression scalars by repeating/broadcasting them into the embedding dimension follows the standard additive formulation widely used across transformers. Because the raw spatial coordinates $(x,y)$ and gene expression values are low-dimensional, a direct addition to the $d$-dimensional positional embeddings requires a map into $\mathbb{R}^{d}$. Repetition provides a simple, parameter-free way to match dimensionality while maintaining the relative scale and directional structure of the original signals. Importantly, we show that initialization that relies purely on an MLP (“No PE” in Table 4). In this formulation, we pass the spatial co-ordinates and gene expression into the latent space using an MLP. This leads to a marked drop in performance, indicating that the positional encodings play an essential role. In Table A10, we ran an additional pretraining run that combines sinusoidal PEs with an MLP-based mapping of coordinates and expression values.
Table A10: PE initialization ablation.
PEs Charville-Outcome Skin-Imputation SEA-Cell classification
No PE ($\text{MLP}(\mathbf{P})$) $0.523\pm 0.010$ $0.458\pm 0.003$ $0.220\pm 0.034$
Sinusoidal PE ($\text{MLP}(\mathbf{P})+PE_{c}$) $0.857\pm 0.045$ $0.538\pm 0.026$ $\mathbf{0.999\pm 0.001}$
HEIST PE ($PE_{c}+{\mathbf{P}}_{c}^{\text{repeat}}$) $\mathbf{0.861\pm 0.086}$ $\mathbf{0.807\pm 0.020}$ $0.995\pm 0.015$
Batch size ablation. To evaluate how smaller block sizes affect performance, we ran a pre-training run with the HEIST pipeline using smaller block sizes. We see a slight decrease in the results except for cell classification where the results became almost perfect. This is because batches would provide less microenvironmental cues in each step, leading to the slight drop in performance.
Table A11: Batch size ablation
Batch size Charville-Outcome Skin-Imputation SEA-Cell classification
128 $0.783\pm 0.003$ $0.791\pm 0.007$ $\mathbf{0.999\pm 0.001}$
256 $\mathbf{0.861\pm 0.086}$ $\mathbf{0.807\pm 0.020}$ $0.995\pm 0.015$